OpenStudy (anonymous):

Solve the linear equation. 2(x + 6) + 6 = 3(x + 5) + 7 A. {6} B. {-4} C. {12} D. {9} I pick D

OpenStudy (anonymous):

D is not correct

OpenStudy (anonymous):

The first step is to distribute the 2 and the 3 outside the parenthesis. 2x+12+6=3x+15+7 Then add like terms 2x+18=3x+22 Then move the variables/numbers 18=x+22 -4=x

OpenStudy (anonymous):

-4 is correct

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

thank you; I understand now

OpenStudy (anonymous):

thank you both

OpenStudy (anonymous):

any time

OpenStudy (anonymous):

welcome